# Decentralized real-time iterations for distributed NMPC

### v2

@@ 3 $i$ 57 7 GN 2 3 1104 640 320 304
TABLE II: Computation times per NMPC step and closed-loop performance $J_{\text{cl}}$ for a prototypical Matlab implementation. The fast computation times in test cases one and three demonstrate the real-time feasibility of decentralized real-time iterations.
@@
VI-C Parameter estimation and validity of assumptions
The OCP design discussed in subsection VI-A meets the value function requirements (3) in Assumption 1 and the pendulum dynamics satisfy Assumption 4. Numerical a-posteriori analyses further show that the KKT points found by ipopt satisfy the regularity Assumption 6 if the system is close to the setpoint. The continuity conditions in Assumptions 1 and 2 thus hold for the identified local minima. Finally, Assumption 5 is satisfied, because the pendulum dynamics are sufficiently smooth.
The bounds (11) on the ADMM iterations $l_{\mathrm{max}}$ and (29) on $\bar{\delta}$ offer a compromise between optimiality, i.e., large $l_{\mathrm{max}}$, and sampling frequency, i.e., small $\bar{\delta}$. To obtain stability for a large sampling interval $\bar{\delta}$ via (29), the outer contraction factor $a_{p}$ must be small. By Theorem 1, this requires $a$ to be small, which results in more ADMM iterations $l_{\mathrm{max}}$ via (11). That is, for low sampling frequencies, ADMM must solve the SQP subproblems to greater accuracy. To estimate the number of ADMM iterations $l_{\mathrm{max}}$ sufficient to guarantee stability for the sampling interal $\delta=40\,$ms chosen in the simulations, we proceed as follows: We estimate the constants for calculating $\delta_{5}$ in Appendix A via simulations close to the setpoint with the ideal centralized NMPC feedback law $\kappa(x)$ using IPOPT, similarly to [Zanelli2021]. These simulations yield constants $a_{1}=0.5326$, $a_{2}=7.1530$, $a_{3}=0.2113$, $L_{f,x}^{\mathrm{c}}=86.2704$, $L_{f,u}^{\mathrm{c}}=3.6685$, $L_{p,x}=51.3647$, $L_{V,x}=1.7908$, and $L_{V,p}=207.1090$. Then, the QP approximation of the OCP at the setpoint yields $c_{1}=1.7321$ and $c_{2}=251.5737$. Finally, we sample the ADMM LTI dynamics (22) in simulations for random $w$ and obtain $a_{w}=0.9989$. As a result, $l_{\mathrm{max}}=24007$ ADMM iterations are sufficient to guarantee stability for $\delta_{5}=40\,$ms. Compared with the settings in Table I, this estimate is quite conservative. Here, the conservatism primarily stems from the Lipschitz constants in the system dynamics and the contraction factor $a_{w}$. In hardware experiments and simulations, we observe good control performance already for 2–30 ADMM iterations per MPC step, depending on the application [Stomberg2023, Stomberg2024b, Stomberg2024c].
VII Conclusion
@@
Appendix A Centralized RTI Stability
This section summarizes the steps for computing the sufficent sampling interval $\bar{\delta}>0$ and optimizer initialization radius $\tilde{r}_{p}>0$ which guarantee stability of centralized RTI [Zanelli2021]. First, we recall the following two results which yield constants $r_{x},r_{p},\delta_{1}>0$.
Lemma 7 (Optimizer perturbation [Zanelli2021])
Let Assumptions 2 and 3 hold. Then, there exist positive constants $r_{x}\leq\hat{r}_{x}$ and $r_{p}\leq\hat{r}_{p}$ such that $\|p^{k+1}-\bar{p}(x^{\prime})\|\leq a_{p}\|p^{k}-\bar{p}(x)\|+L_{p,x}a_{p}\|x^{\prime}-x\|$ for all $x\in\mathbb{X}_{\bar{V}}$, $p\in\mathcal{B}(\bar{p}(x),r_{z})$, and $x^{\prime}\in\mathcal{B}(x,r_{x})$. $\square$
Lemma 8 (Lipschitz discrete-time dynamics [Zanelli2021])
Let Assumption 4 hold. Then, there exists a positive finite constant $\delta_{1}$ such that, for all $x\in\mathbb{X}_{\bar{V}}$, $p\in\mathcal{B}(\bar{p}(x),r_{p})$, and $\delta\leq\delta_{1}$,
@@ Appendix A Centralized RTI Stability
where $L_{f,x}^{\delta_{1}}\doteq e^{L_{f,x}^{\mathrm{c}}\delta_{1}}L_{f,x}^{\mathrm{c}}$ and $L_{f,u}^{\delta_{1}}\doteq e^{L_{f,x}^{\mathrm{c}}\delta_{1}}L_{f,u}^{\mathrm{c}}$. $\square$
@@
$\delta_{3}\doteq\min\left\{\delta_{0},\delta_{1},\frac{r_{x}}{\eta r_{\bar{V}}+L_{f,u}^{\delta_{1}}r_{p}},\frac{r_{p}(1-a_{p})}{L_{p,x}a_{p}\left(L_{f,u}^{\delta_{1}}r_{p}+\eta r_{\bar{V}}\right)}\right\},$
$\kappa\doteq a_{p}\left(1+\delta_{3}L_{p,x}L_{f,u}^{\delta_{1}}\right)$, $L_{V}\doteq 2\bar{V}^{1/2}L_{V,x}$, ${\bar{a}\doteq a_{3}/a_{2}}$, $L_{e}\doteq L_{V}L_{f,u}^{\delta_{1}}$, $L_{V,p}\doteq L_{f,u}^{\mathrm{c}}e^{\delta_{1}L_{f,x}^{\mathrm{c}}}L_{V,x}$, ${\beta\doteq(\bar{a}\sqrt{a_{1}})/(4L_{p,x}a_{p}\eta)}$,
$\delta_{4}^{\prime}\doteq\frac{(1-\kappa)\tilde{r}_{p}\sqrt{a_{1}}}{\bar{V}^{1/2}L_{p,x}a_{p}\eta},\quad\text{and}\quad\delta_{5}\doteq\frac{\beta(1-\kappa)}{L_{V,p}}.$
The sampling interval $\bar{\delta}$ and optimizer initialization radius $\tilde{r}_{p}$ sufficient for closed-loop stability read [Zanelli2021]
$\bar{\delta}\doteq\min\{\delta_{3},\delta_{4}^{\prime},\delta_{5}\}\quad\text{and}\quad\tilde{r}_{p}\doteq\min\left\{r_{p},\bar{a}\bar{V}/L_{e}\right\}.$ (29)
Appendix B NLP Formulation and Decentralized ADMM
@@
Example 1 (OCP as partially separable NLP [Stomberg2022])
@@
Example 1 reformulates coupled dynamics. Coupled costs (1a) and constraints (1e) can be treated similarly.
The reformulation allows to decentralize ADMM as follows. The sparsity in $E_{1}$ and $E_{2}$ together with $c=0$ allow to view NLP (5) as a so-called consensus problem, where each subsystem optimizes over a selection of the centralized decision variables $(\bar{\boldsymbol{x}},\bar{\boldsymbol{u}})$. Specifically for Example 1, Subsystem 1 optimizes over $(\bar{x}_{1}(0),\bar{x}_{1}(1),\bar{u}_{1}(0))$ and Subsystem 2 optimizes over $(\bar{x}_{1}(0),\bar{x}_{2}(0),\bar{x}_{2}(1))$. A derivation of the decentralized implementation of Step (8b) for general consensus problems is provided in [Boyd2011, Ch. 7.2]. To make a closer connection to NLP (5), recall from the proof of Lemma 4 that the $z$-update in ADMM is given by $z^{l+1}=M_{\mathrm{avg}}(y^{l}+\gamma^{l}/\rho).$ For Example 1, the ADMM averaging matrix $M_{\mathrm{avg}}$ reads
@@
$[z_{1}]_{1}^{l+1}=[z_{2}]_{3}^{l+1}=\frac{[y_{1}]_{1}^{l+1}+[y_{2}]_{3}^{l+1}+([\gamma_{1}]_{1}^{l}+[\gamma_{2}]_{3}^{l})/\rho}{2}.$
A common approach to decentralize ADMM in DMPC is to send the variables on the right hand side of to the above average to the subsystem with the original state, i.e., Subsystem 1 in Example 1. Then, the average is evaluated and the result is sent to the subsystems with the copied states, i.e., Subsystem 2 in Example 1. For more details, see [Bestler2019, Stomberg2023, Stomberg2024b].
[…]

### v3

@@ 3 $i$ 57 7 GN 2 3 1104 640 320 304
Table 2: Computation times per NMPC step and closed-loop performance $J_{\text{cl}}$ for a prototypical Matlab implementation. The fast computation times in test cases one and three demonstrate the real-time feasibility of decentralized real-time iterations.
@@
6.3 Parameter Estimation and Validity of Assumptions
The OCP design discussed in subsection 6.1 meets the value function requirements (3) in Assumption 1 and the pendulum dynamics satisfy Assumption 4. Numerical a-posteriori analyses further show that the KKT points found by IPOPT satisfy the regularity Assumption 6 if the system is close to the setpoint. The continuity conditions in Assumptions 1 and 2 thus hold for the identified local minima. Finally, Assumption 5 is satisfied, because the pendulum dynamics are sufficiently smooth.
The bounds (11) on the ADMM iterations $l_{\mathrm{max}}$ and (31) on $\bar{\delta}$ offer a compromise between optimiality, i.e., large $l_{\mathrm{max}}$, and sampling frequency, i.e., small $\bar{\delta}$. To obtain stability for a large sampling interval $\bar{\delta}$ via (31), the outer contraction factor $a_{p}$ must be small. By Theorem 10, this requires $a$ to be small, which results in more ADMM iterations $l_{\mathrm{max}}$ via (11). That is, for low sampling frequencies, ADMM must solve the SQP subproblems to greater accuracy. To estimate the number of ADMM iterations $l_{\mathrm{max}}$ sufficient to guarantee stability for the sampling interal $\delta=40\,$ms chosen in the simulations, we proceed as follows: We estimate the constants for calculating $\delta_{5}$ in Appendix 8 via simulations close to the setpoint with the ideal centralized NMPC feedback law $\kappa_{\mathrm{c}}(x)$ using IPOPT, similarly to [7]. These simulations yield constants $a_{1}=0.5326$, $a_{2}=7.1530$, $a_{3}=0.2113$, $L_{f,x}^{\mathrm{c}}=86.2704$, $L_{f,u}^{\mathrm{c}}=3.6685$, $L_{p,x}=51.3647$, $L_{V,x}=1.7908$, and $L_{V,p}=207.1090$. Then, the QP approximation of the OCP at the setpoint yields $c_{1}=1.7321$ and $c_{2}=251.5737$. Finally, we sample the ADMM LTI dynamics (22) in simulations for random $w$ and obtain $a_{w}=0.9989$. As a result, $l_{\mathrm{max}}=24007$ ADMM iterations are sufficient to guarantee stability for $\delta_{5}=40\,$ms. Compared with the settings in Table 1, this estimate is quite conservative. Here, the conservatism primarily stems from the Lipschitz constants in the system dynamics and the contraction factor $a_{w}$. The conservatism could thus be reduced by more accurate estimates for these constants, for instance through tighter ADMM convergence guarantees or elaborate simulation. In hardware experiments and simulations, we observe good control performance already for 2–30 ADMM iterations per MPC step, depending on the application [29, 30, 31].
7 Conclusion
@@
[…]
8 Centralized RTI Stability
This section summarizes the steps for computing the sufficent sampling interval $\bar{\delta}>0$ and optimizer initialization radius $\tilde{r}_{p}>0$ which guarantee stability of centralized RTIs [7].
Lemma 20 (Lipschitz discrete-time dynamics [7]).
Let Assumption 4 hold. Then, there exists a positive finite constant $\delta_{1}$ such that, if $x\in\mathbb{X}_{\bar{V}}$, $p\in\mathcal{B}(\bar{p}(x),r_{p}^{\prime})$, and $\delta\leq\delta_{1}$, then
@@ 8 Centralized RTI Stability
where $L_{f,x}^{\delta_{1}}\doteq e^{L_{f,x}^{\mathrm{c}}\delta_{1}}L_{f,x}^{\mathrm{c}}$ and $L_{f,u}^{\delta_{1}}\doteq e^{L_{f,x}^{\mathrm{c}}\delta_{1}}L_{f,u}^{\mathrm{c}}$. Moreover, if $\delta\leq\delta_{1}$, then
$\left\|f^{\delta}(x,u^{\prime})-f^{\delta}(x,u)\right\|\leq\delta L_{f,u}^{\delta_{1}}\left\|u^{\prime}-u\right\|$ (29)
for all $x\in\mathbb{X}_{\bar{V}}$, all $u^{\prime}=M_{u,p}p^{\prime}$, $u=M_{u,p}p$ such that $p,p^{\prime}\in\mathcal{B}(\bar{p}(x),r_{p}^{\prime})$. $\square$
Lemma 21 (Forward invariance under perturbed optimizer).
Let Assumptions 1 and 4 hold. Then, there exists a positive constant $r_{p}^{\prime\prime}\leq r_{p}^{\prime}$ such that $f^{\delta}(x,M_{u,p}p)\in\mathbb{X}_{\bar{V}}$ for all $x\in\mathbb{X}_{\bar{V}}$, all $p\in\mathcal{B}(\bar{p}(x),r_{p}^{\prime\prime})$, and if $\delta\leq\delta_{1}$. $\square$
Proof 8.1.
We define the shorthands $x_{+}^{\star}\doteq f^{\delta}(x,M_{u,p}\bar{p}(x))$ and $x_{+}\doteq f^{\delta}(x,M_{u,p}{p})$. Let $r_{p}^{\prime\prime}\leq r_{p}^{\prime}$ with $r_{p}^{\prime}$ from Assumption 4. By the Lyapunov decrease condition in Assumption 1, we have that $V(x_{+}^{\star})\leq V(x)-\delta a_{3}\|x\|^{2}$. If $x=0$, then $V(x)=0$ and thus $V(x_{+}^{\star})=0<\bar{V}$. If $x\neq 0$, then $V(x_{+}^{\star})\leq\bar{V}-\delta a_{3}\|x\|^{2}<\bar{V}$. Thus, there exists a constant $\epsilon_{V}>0$ such that $V(x_{+}^{\star})+\epsilon_{V}\leq\bar{V}$. Recall that, by Assumption 1, $\mathbb{X}_{\bar{V}}$ lies in the interior of $\mathbb{X}_{0}$ and that $V$ is continuous over $\mathbb{X}_{0}$. Thus, there exists a constant $\epsilon_{x}>0$ such that $|V(x^{\prime})-V(x_{+}^{\star})|<\epsilon_{V}$ for all $\|x^{\prime}-x_{+}^{\star}\|<\varepsilon_{x}$ [42, App. A.11]. By inserting $\|p-\bar{p}(x)\|\leq r_{p}^{\prime\prime}$ into (29), we have that $\|x_{+}-x_{+}^{\star}\|\leq\delta L_{f,u}^{\delta_{1}}r_{p}^{\prime\prime}$. For $r_{p}^{\prime\prime}>0$ sufficiently small, $\|x_{+}-x_{+}^{\star}\|<\epsilon_{x}$ and so $V(x_{+})<V(x_{+}^{\star})+\epsilon_{V}\leq\bar{V}$. By the definition of the level set $\mathbb{X}_{\bar{V}}$, we obtain that $x_{+}\in\mathbb{X}_{\bar{V}}$, which concludes the proof.
Define the shorthands $x_{t}\doteq x(t)$, $p_{t}\doteq p^{k_{\mathrm{max}}}(t)$, and $p^{\star}_{t}\doteq\bar{p}(x(t))$, and similarly for $\cdot_{t+1}$. We next adapt [7, Lem. 11] to the q-linear optimizer convergence Assumption 3.
Lemma 22 (Contraction).
Let Assumptions 14 hold and consider the system-optimizer dynamics (4). Then, there exist positive constants $r_{p}\leq\min\{\hat{r}_{p},r_{p}^{\prime\prime}\}$ and $r_{x}\leq\hat{r}_{x}$ such that, for all $x_{t}\in\mathbb{X}_{\bar{V}}$ and all $\|p_{t}-p^{\star}_{t}\|\leq r_{p}$, the following holds. If $\delta\leq\delta_{1}$ and if $\|x_{t+1}-x_{t}\|\leq r_{x}$, then
$\|p_{t+1}-p^{\star}_{t+1}\|\leq a_{p}\|p_{t}-p^{\star}_{t}\|+a_{p}L_{p,x}\|x_{t+1}-x_{t}\|.$ (30)
$\square$
Proof 8.2.
Since $r_{p}\leq r_{p}^{\prime\prime}$, Lemma 21 yields that $x_{t+1}=f^{\delta}(x_{t},M_{u,p}p_{t})\in\mathbb{X}_{\bar{V}}$. Assumption 2 on the Lipschitz continuity of the KKT point thus states that
$\|p^{\star}_{t+1}-p^{\star}_{t}\|\leq L_{p,x}\|x_{t+1}-x_{t}\|\leq L_{p,x}r_{x}.$
By further setting $r_{p}\leq\min\{r_{p}^{\prime\prime},\hat{r}_{p}-L_{p,x}r_{x}\}$ with $r_{x}<\min\{\hat{r}_{x},\hat{r}_{p}/L_{p,x}\}$}, we obtain
$\|p_{t}-p^{\star}_{t+1}\|$ $\leq\|p_{t}-p^{\star}_{t}\|+\|p^{\star}_{t+1}-p^{\star}_{t}\|$
$\leq r_{p}+L_{p,x}r_{x}\leq\hat{r}_{p}.$
Thus, $\|p_{t}-p^{\star}_{t+1}\|$ lies inside the q-linear convergence region of the optimizer. Assumptions 2 and 3 hence yield
$\|p_{t+1}-p^{\star}_{t+1}\|$ $\leq(a_{p})^{k_{\mathrm{max}}}\|p_{t}-p_{t+1}^{\star}\|$
$\leq a_{p}\|p_{t}-p_{t+1}^{\star}\|$
$\leq a_{p}\|p_{t}-p_{t}^{\star}\|+a_{p}\|p^{\star}_{t+1}-p^{\star}_{t}\|$
$\leq a_{p}\|p_{t}-p_{t}^{\star}\|+a_{p}L_{p,x}\|x_{t+1}-x_{t}\|,$
where we have used that $(a_{p})^{k_{\mathrm{max}}}\leq a_{p}<1$. This finishes the proof.
@@
$\delta_{3}\doteq\min\left\{\delta,\delta_{1},\frac{r_{x}}{\eta r_{\bar{V}}+L_{f,u}^{\delta_{1}}r_{p}},\frac{r_{p}(1-a_{p})}{L_{p,x}a_{p}\left(L_{f,u}^{\delta_{1}}r_{p}+\eta r_{\bar{V}}\right)}\right\},$
$\kappa\hskip-1.42262pt\doteq\hskip-1.42262pta_{p}\left(1\hskip-1.42262pt+\hskip-1.42262pt\delta_{3}L_{p,x}L_{f,u}^{\delta_{1}}\right)$, $L_{V}\doteq 2\bar{V}^{1/2}L_{V,x}$, $\bar{a}\doteq a_{3}/a_{2}$}, $L_{e}\doteq L_{V}L_{f,u}^{\delta_{1}}$, $L_{V,p}\doteq L_{f,u}^{\mathrm{c}}e^{\delta_{1}L_{f,x}^{\mathrm{c}}}L_{V,x}$, and $\beta^{\prime}\doteq(\bar{a}\sqrt{a_{1}})/(4L_{p,x}a_{p}\eta)$}.
The sampling interval $\bar{\delta}$ and optimizer initialization radius $\tilde{r}_{p}$ sufficient for closed-loop stability read [7]
$\bar{\delta}\doteq\min\{\delta_{3},\delta_{4}^{\prime},\delta_{5}\}\quad\text{and}\quad\tilde{r}_{p}\doteq\min\left\{r_{p},\bar{a}\bar{V}/L_{e}\right\},$ (31)
where
$\delta_{4}^{\prime}\doteq\frac{(1-\kappa)\tilde{r}_{p}\sqrt{a_{1}}}{\bar{V}^{1/2}L_{p,x}a_{p}\eta},\quad\text{and}\quad\delta_{5}\doteq\frac{\beta^{\prime}(1-\kappa)}{L_{V,p}}.$
9 NLP Formulation and Decentralized ADMM
@@
Example 9.1 (OCP as partially separable NLP [20]).
@@
Example 9.1 reformulates coupled dynamics. Coupled costs (1a) and constraints (1e) can be treated similarly. The reformulation allows to decentralize ADMM as follows. The sparsity in $E_{1}$ and $E_{2}$ together with $c=0$ allow to view NLP (5) as a so-called consensus problem, where each subsystem optimizes over a selection of the centralized decision variables $(\bar{\boldsymbol{x}},\bar{\boldsymbol{u}})$. Specifically for Example 9.1, Subsystem 1 optimizes over $(\bar{x}_{1}(0),\bar{x}_{1}(1),\bar{u}_{1}(0))$ and Subsystem 2 optimizes over $(\bar{x}_{1}(0),\bar{x}_{2}(0),\bar{x}_{2}(1))$. A derivation of the decentralized implementation of Step (8b) for general consensus problems is provided in [34, Ch. 7.2]. To make a closer connection to NLP (5), recall from the proof of Lemma 6 that the $z$-update in ADMM is given by $z^{l+1}=M_{\mathrm{avg}}(y^{l}+\gamma^{l}/\rho).$ For Example 9.1, the ADMM averaging matrix $M_{\mathrm{avg}}$ reads
@@
$z_{1}]_{1}^{l+1}=[z_{2}]_{3}^{l+1}=\frac{[y_{1}]_{1}^{l+1}+[y_{2}]_{3}^{l+1}+([\gamma_{1}]_{1}^{l}+[\gamma_{2}]_{3}^{l})/\rho}{2}.$
A common approach to decentralize ADMM in DMPC is to send the variables on the right hand side of the above average to the subsystem with the original state, i.e., Subsystem 1 in Example 9.1. Then, the average is evaluated and the result is sent to the subsystems with the copied states, i.e., Subsystem 2 in Example 9.1. For more details, see [16, 29, 30].